# A Novel Discrete-time Model of Information Diffusion on Social Networks Considering Users Behavior

### v1

@@ Abstract
In this paper, we introduce the SDIR (Susceptible–Delayable–Infected–Recovered) model, an extension of the classical SIR epidemic framework, to provide a more explicit characterization of user behavior in online social networks. The newly merged state D (delayable) represents users who have received the information but delayed its spreading and may eventually choose not to share it at all. Based on the mean-field approximation method, we derive the dynamical equations of the model and investigate its convergence and stability conditions. Under these conditions, we further propose a Sandwich approximation algorithm for the edge-deletion problem, aiming to minimize the influence of information diffusion by identifying approximate solutions.
@@ I Introduction
In this paper, we present a novel discrete-time SDIR model, a new extension of the classical discrete SIR framework. The SDIR model incorporates an intermediate state, D (Delayable), between S and I. This new state captures scenarios where a node that receives information does not immediately become “infected,” but may instead exhibit a delay before spreading the information-or may ultimately choose not to spread it at all. This behavioral shade reflects realistic interactions in online social networks, where users exercise discretion in processing, accepting, and sharing information. Each seed node corresponds to a social media account (e.g., a Facebook/Twitter/YouTube/TikTok user account or fan page) that initiates diffusion to its followers at the beginning. In reality, many users either ignore their received information or share it only after a delay in judgment. By introducing the Delayable state, the SDIR model offers a more accurate representation of individual-level behavior in information diffusion.
@@ II Proposed Model
$D_{i}(t+1)=S_{i}(t)(1-\alpha_{i}(t))$
$\Biggl(1-\prod_{j=1}^{n}\bigl(1-\beta_{ij}(t)I_{j}(t)\bigr)\Biggr)$
@@
$R_{i}(t+1)=\delta_{i}(t)I_{i}(t)+\delta^{\prime}_{i}(t)I^{\prime}_{i}(t)+R_{i}(t)$ (3)
@@ III Sufficient Condition for Global Convergence
One difference between the SDIR model and the SEIR model is that the delayed state in the SDIR model is still considered an infected state, while the SEIR model only considers E as an exposed state. Therefore, when considering global stability in the SDIR model, we need to find a sufficient condition for $\bm{x}(t)$ and $\bm{y}(t)$ to both approach the zero state. Consider a vector $\bm{q}\in(0,1]^{n}$ which has $n$ elements $q_{1}$, $q_{2}$$\cdots,$ $q_{n}$, then define
@@ IV-A The main problem
Let the vector $\bm{m}(t)\coloneqq\bm{x}(t)+\bm{y}(t)+\bm{r}(t)$ and $\bm{m}^{\star}\in\mathbb{R}^{n}$ consisting of elements $m^{\star}_{i}=\sup_{t}{m_{i}(t)}$.
@@
The quantity $\|\bm{m}^{*}-\bm{m}(0)\|_{1}$ is called the estimated infection amount on $G$ after the diffusion process ends.
@@
Given a directed graph $G(V,E)$ with $|V|=n$, representing SDIR diffusion model, an initial state vector $\bm{x}(0)\in[0,1]^{n}$ and $\bm{y}(0)$, $\bm{r}(0)$ such that $\bm{x}(0)+\bm{y}(0)+\bm{r}(0)\in[0,1]^{n}$. Let $Q$ be a candidate set of edges such that $Q\subseteq E$ and a positive integer $k$ satisfying $k\leq|Q|$. Find a set of edges $P\subseteq Q,|P|\leq k$ to delete from the graph such that the infection amount on $G_{-P}$ is minimized i.e., find
@@ IV-C Supermodular Lower Bound
One noteworthy point in the SDIR model is that one can find a lower bound function that is also supermodular. Hence, instead of running the greedy algorithm only for the upper bound function $\sigma^{U}(.)$, one can run the greedy algorithm once more for the following lower bound function $\sigma^{L}(.)$. This is precisely the idea of the Sandwich approximation and will be presented in more detail in the next section. Let $\bm{N}_{-P}\coloneqq\bm{I}-\bm{D}+\bm{A}\bm{S}(0)\bm{B}_{-P}$.
@@
$\sigma^{U}(\varnothing)-\sigma(P_{\text{Sand}})\geq$
$\max\left\{\frac{\sigma^{U}(\varnothing)-\sigma(P_{L})}{\sigma^{U}(\varnothing)-\sigma^{L}(P_{L})},\frac{\sigma^{U}(\varnothing)-\sigma^{U}(P^{\star})}{\sigma^{U}(\varnothing)-\sigma(P^{\star})}\right\}$
$(1-1/e-\epsilon)(\sigma^{U}(\varnothing)-\sigma(P^{\star})),$
@@ V Algorithms
Initialize $P\leftarrow\varnothing$
@@
VI Conclusion
We studied the discrete-time SDIR model, a new modified SIR model. We showed that the sufficient condition for the convergence and the stability of the SDIR model. Furthermore, we investigate the problem that minimize the number of infections where deleting some edges. Two algorithms, greedy and Sandwich algorithms were proposed to solve the minimizing infections problem.
@@ VI Conclusion
Therefore, $\bm{x}(t)+\bm{Q}\bm{y}(t)\leq\bm{M}^{t}(\bm{x}(0)+\bm{Q}\bm{y}(0))$. If $\rho(\bm{M})$ is upper-bounded by $1-\epsilon$, then $\|\bm{x}(t)+\bm{Q}\bm{y}(t)\|\leq(1-\epsilon)^{t}\|\bm{x}(0)+\bm{Q}\bm{y}(0)\|$, which forces $\bm{x}(t)$ and $\bm{y}(t)$ to converge to 0 as $t\to\infty$. This completes the proof.
VI-B Proof of Theorem IV.3
@@
Returning to Theorem IV.5. We have $\bm{x}(t+1)=\bm{N}_{-P}\bm{x}(t)+\bm{W}\bm{y}(t)=\bm{N}_{-P}^{\,t+1}\bm{x}(0)+\sum_{l=0}^{t}\bm{N}_{-P}^{\,t-l}\bm{W}\bm{y}(l),\forall t\geq 0$. It is easy to see that $\bm{y}(t)\geq\bm{F}\bm{y}(t-1)$ with $\bm{F}=\bm{I}-\bm{W}-\bm{D}^{\prime}$, it follows that $\bm{y}(t)\geq\bm{F}^{t}\bm{y}(0)$ for all $t\geq 0$.
$\Rightarrow\bm{m}(t+1)-\bm{m}(0)=\bm{S}(0)\bm{B}_{-P}\sum_{s=0}^{t}\bm{x}(s)$
$=\bm{S}(0)\bm{B}_{-P}\Bigg(\sum_{s=1}^{t}\Big(\bm{N}_{-P}^{s}\bm{x}(0)$
@@
$\bm{S}(0)\bm{B}_{-P}\sum_{l=0}^{t}\bm{N}_{-P}^{\,l}\bm{x}(0)$
$+\bm{S}(0)\bm{B}_{-P}\sum_{s=1}^{t}\sum_{l=0}^{s-1}\bm{N}_{-P}^{\,s-1-l}\bm{W}\bm{F}^{\,l}\bm{y}(0)$
$=\bm{S}(0)\bm{B}_{-P}\sum_{l=0}^{t}\bm{N}_{-P}^{\,l}\bm{x}(0)$
$+\bm{S}(0)\bm{B}_{-P}\sum_{s=0}^{t-1}\sum_{l=0}^{s}\bm{N}_{-P}^{\,s-l}\bm{F}^{\,l}\bm{W}\bm{y}(0),$
for all $t\geq 1$.
Letting $t\rightarrow\infty$ and applying Lemma VI.1 yields $\bm{m}^{\star}-\bm{m}(0)\geq\bm{S}(0)\bm{B}_{-P}\Big((\bm{I}-\bm{N}_{-P})^{-1}\bm{x}(0)+(\bm{I}-\bm{N}_{-P})^{-1}(\bm{I}-\bm{F})^{-1}\bm{W}\bm{y}(0)\Big)$. Equivalently, $\|\bm{m}^{\star}-\bm{m}(0)\|_{1}\geq\bm{1}^{\top}\bm{A}^{-1}\big(\bm{D}(\bm{I}-\bm{N}_{-P})^{-1}-\bm{I}\big)\big(\bm{x}(0)+\bm{W}(\bm{W}+\bm{D}^{\prime})^{-1}\bm{y}(0)\big)$.

### v2

@@ Abstract
In this paper, we introduce the SDIR (Susceptible–Delayable–Infected–Recovered) model, an extension of the classical SIR epidemic framework, to provide a more explicit characterization of user behavior in online social networks. The newly merged state D (delayable) represents users who have received the information but delayed its spreading and may eventually choose not to share it at all. Based on the mean-field approximation method, we derive the dynamical equations of the model and investigate its convergence and stability conditions. Under these conditions, we further propose a greedy algorithm and a sandwich approximation algorithm for the edge-deletion problem, aiming to minimize the influence of information diffusion by identifying approximate solutions.
@@ I Introduction
In this paper, we present a novel discrete-time SDIR model, a new extension of the classical discrete SIR framework. The SDIR model incorporates an intermediate state, D (Delayable), between S and I. This new state captures scenarios where a node that receives information does not immediately become “infected,” but may instead exhibit a delay before spreading the information-or may ultimately choose not to spread it at all. This behavioral shade reflects realistic interactions in online social networks, where users exercise discretion in processing, accepting, and sharing information. Each seed node corresponds to a social media account (e.g., a Facebook/Twitter/YouTube/TikTok user account or fan page) that initiates diffusion to its followers at the beginning. In reality, many users either ignore their received information or share it only after a delay in judgment. User procrastination on online social networks has been examined as an inherent behavioral factor in the use of such platforms [6, 25]. By introducing the Delayable state, the SDIR model offers a more accurate representation of individual-level behavior in information diffusion.
@@ II Proposed Model
$D_{i}(t+1)=$ $S_{i}(t)(1-\alpha_{i}(t))\Biggl(1-\prod_{j=1}^{n}\bigl(1-\beta_{ij}(t)I_{j}(t)\bigr)\Biggr)$
@@
$R_{i}(t+1)=$ $\delta_{i}(t)I_{i}(t)+\delta^{\prime}_{i}(t)D_{i}(t)+R_{i}(t)$ (3)
@@ III Sufficient Condition for Global Convergence
One difference between the SDIR model and the SEIR model in [21] is that the delayed state in the SDIR model is still considered an infected state, while the SEIR model only considers E as an exposed state. Therefore, when considering global stability in the SDIR model, we need to find a sufficient condition for $\bm{x}(t)$ and $\bm{y}(t)$ to both approach the zero state. Consider a vector $\bm{q}\in(0,1]^{n}$ which has $n$ elements $q_{1}$, $q_{2}$$\cdots,$ $q_{n}$, then define
@@ IV-A The main problem
Let the vector $\bm{m}(t)\coloneqq\bm{x}(t)+\bm{y}(t)+\bm{r}(t)$ and $\bm{m}^{\star}\in\mathbb{R}^{n}$ consisting of elements $m^{\star}_{i}=\sup\limits_{t}{m_{i}(t)}$.
@@
The quantity $\|\bm{m}^{*}-\bm{m}(0)\|_{1}$ is called the number of increased infections on $G$ after the diffusion process ends.
@@
Given a directed graph $G(V,E)$ with $|V|=n$, representing SDIR diffusion model, an initial state vector $\bm{x}(0)\in[0,1]^{n}$ and $\bm{y}(0)$, $\bm{r}(0)$ such that $\bm{x}(0)+\bm{y}(0)+\bm{r}(0)\in[0,1]^{n}$. Let $Q$ be a candidate set of edges such that $Q\subseteq E$ and a positive integer $k$ satisfying $k\leq|Q|=q$. Find a set of edges $P\subseteq Q,|P|\leq k$ to delete from the graph such that the infection amount on $G_{-P}$ is minimized i.e., find
@@ IV-C Supermodular Lower Bound
A noteworthy point in the SDIR model is that one can find a lower bound function that is also supermodular. Hence, instead of running the greedy algorithm only for the upper bound function $\sigma^{U}(.)$, one can run the greedy algorithm once more for the following lower bound function $\sigma^{L}(.)$. This is precisely the idea of the Sandwich approximation and will be presented in more detail in the next section. Let $\bm{N}_{-P}\coloneqq\bm{I}-\bm{D}+\bm{A}\bm{S}(0)\bm{B}_{-P}$.
@@
$\sigma^{U}(\emptyset)-\sigma(P_{\text{Sand}})\geq$ $\max{\left\{\frac{\sigma^{U}(\emptyset)-\sigma(P_{L})}{\sigma^{U}(\emptyset)-\sigma^{L}(P_{L})},\frac{\sigma^{U}(\emptyset)-\sigma^{U}(P^{\star})}{\sigma^{U}(\emptyset)-\sigma(P^{\star})}\right\}}$
$\left(1-\frac{1}{e}-\epsilon\right)(\sigma^{U}(\emptyset)-\sigma(P^{\star})),$
@@ V Algorithms
Initialize $P\leftarrow\emptyset$
@@
VI Experiments
In this section, we present numerical examples to demonstrate the convergence and effectiveness of the proposed algorithms. We evaluate the performance of our method on a real-world dataset and compare it against some known strategies.
VI-A Comparison with Heuristic Algorithms
First, we compare the proposed SAA with two heuristic algorithms, namely Max-degree [2] and Random [7], applied to the SDIR model. The algorithms are defined as follows:
  1. 1.
    Max-degree: This algorithm iteratively removes the edge incident to the node with the highest weighted degree at each step.
  2. 2.
    Random: This algorithm removes an edge selected uniformly at random at each iteration. Note that the procedure for determining $P_{0}$ is integrated into this selection process.
We conducted experiments using the real-world contact network dataset collected in Haslemere, England [14]. The parameter settings and preprocessing steps follow the methodology described in Yi et al [34]. The new simulation parameters are initialized as follows: The initial state $\bm{y}(0)$ is set to zero for all nodes, except for five seed nodes which are assigned random values in the interval $[0,0.05]$. Also, the initial state $\bm{x}(0)$ is set to zero for all nodes, except for five seed nodes which are assigned random values in the interval $[0.8,0.85]$. Next, the infection rates $A_{i}$ are distributed based on node categories: the five seed nodes have $A_{i}\in[0.9,1]$; forty-five randomly selected nodes have $A_{i}\in[0.65,0.87]$; and the remaining nodes have $A_{i}\in[0.15,0.35]$. For each node $i\in V$, the weight $W_{i}$ is chosen uniformly at random from $[0.15,0.35]$. To ensure Assumption 1 is satisfied, $D^{\prime}_{i}$ is selected from the interval $\max(0,D_{i}-W_{i}),0.95-W_{i}]$.
Figure 2: Haslemere Network with preprocessing
The simulation results are presented in Fig. 2. It is evident that the solutions obtained from the proposed sandwich bounding functions significantly outperform the two heuristic algorithms. The number of increased infections decreases rapidly within the removal of the first 50 edges. Notably, the upper and lower bounds derived from the sandwich algorithms are very close to each other, indicating the tightness of the approximation with respect to the objective function. It is similar to the results or the complexity of the Greedy algorithm in [34], in our simulations, the running time of our Algorithm 1 for the SDIR model is $O(n^{3}+k(n^{2}+qn))$.
VI-B Convergence Analysis
Figure 3: Convergence of infection states over time
Next, utilizing the same parameter setup, we examine the convergence of the infection states in both the proposed SDIR model and the linear SIR model. The total amount of infection state of the SDIR model is defined as $\bm{x}(t)+\bm{y}(t)||_{1}$. For comparison, the total amount of infection state of the linear SIR model is calculated as $\hat{\bm{x}}(t)||_{1}$, where the initial condition is set to $\hat{\bm{x}}(0)=\bm{x}(0)+\bm{y}(0)$.
The results illustrated in Fig. 3 demonstrate that the SDIR model exhibits faster convergence compared to the linear SIR model. Specifically, the infection states in the SDIR model rapidly decays to $\bm{0}_{n\times 1}$ within the first 10 time steps. In contrast, the linear SIR model exhibits a slower decay, requiring approximately 5 additional time steps to reach a similar extinction state. This empirical observation reinforces the theoretical validity of Theorem III.1 under Assumption 1.
VII Conclusion
We studied the discrete-time SDIR model, a new modified SIR model. We showed that the sufficient condition for the convergence and the stability of the SDIR model. Furthermore, we investigate the problem that minimize the number of infections where deleting some edges. Two algorithms, greedy and Sandwich algorithms were proposed to solve the minimizing infections problem. Some experimental results were given.
@@ VII Conclusion
Therefore, $\bm{x}(t)+\bm{Q}\bm{y}(t)\leq\bm{M}^{t}(\bm{x}(0)+\bm{Q}\bm{y}(0))$. If $\rho(\bm{M})\leq 1-\epsilon$, then $\|\bm{x}(t)+\bm{Q}\bm{y}(t)\|\leq(1-\epsilon)^{t}\|\bm{x}(0)+\bm{Q}\bm{y}(0)\|$, which forces $\bm{x}(t)$ and $\bm{y}(t)$ to converge to 0 as $t\to\infty$. This completes the proof.
VII-B Proof of Theorem IV.3
@@
Returning to Theorem IV.5. We have $\bm{x}(t+1)=\bm{N}_{-P}\bm{x}(t)+\bm{W}\bm{y}(t)=\bm{N}_{-P}^{\,t+1}\bm{x}(0)+\sum_{l=0}^{t}\bm{N}_{-P}^{\,t-l}\bm{W}\bm{y}(l),\forall t\geq 0$. It is easy to see that $\bm{y}(t)\geq\bm{F}\bm{y}(t-1)$ with $\bm{F}=\bm{I}-\bm{W}-\bm{D}^{\prime}$, it follows that $\bm{y}(t)\geq\bm{F}^{t}\bm{y}(0)$ for all $t\geq 0$. So
$\bm{m}(t+1)-\bm{m}(0)=\bm{S}(0)\bm{B}_{-P}\Bigg(\sum_{s=1}^{t}\Big(\bm{N}_{-P}^{s}\bm{x}(0)$
@@
$\bm{S}(0)\bm{B}_{-P}\sum_{l=0}^{t}\bm{N}_{-P}^{\,l}\bm{x}(0)+\bm{S}(0)\bm{B}_{-P}\sum_{s=0}^{t-1}\sum_{l=0}^{s}\bm{N}_{-P}^{\,s-l}\bm{F}^{\,l}\bm{W}\bm{y}(0).$
Setting $t\rightarrow\infty$ and applying Lemma VII.1 yields $\bm{m}^{\star}-\bm{m}(0)\geq\bm{S}(0)\bm{B}_{-P}\Big((\bm{I}-\bm{N}_{-P})^{-1}\bm{x}(0)+(\bm{I}-\bm{N}_{-P})^{-1}(\bm{I}-\bm{F})^{-1}\bm{W}\bm{y}(0)\Big)$. Then, equivalently, $\|\bm{m}^{\star}-\bm{m}(0)\|_{1}\geq\bm{1}^{\top}\bm{A}^{-1}\big(\bm{D}(\bm{I}-\bm{N}_{-P})^{-1}-\bm{I}\big)\big(\bm{x}(0)+\bm{W}(\bm{W}+\bm{D}^{\prime})^{-1}\bm{y}(0)\big)$.